OpenStudy (anonymous):

The ratio of apples to oranges at a fruit stand is 4 to 5. There are 60 oranges at the stand. How many apples are at the fruit stand?

OpenStudy (anonymous):

48 apples

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

4/5 =60 you can do it either 60/4 and 60/5 or 4*60=240/5= 48 apples